OpenStudy (anonymous):

sqrt (2x -3= 5 ) Solve for x?

OpenStudy (anonymous):

\[\sqrt{2x-3}=5\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so, square both sides and get \[2x-3=25\] then solve that one in two steps

OpenStudy (anonymous):

I just basically need to understand how to do these kind of questions

OpenStudy (anonymous):

add 3, divide by 2

OpenStudy (anonymous):

if you have a square root, you need to get rid of it by squaring both sides of the equation

OpenStudy (anonymous):

okay and when you square it, it just stays the same. correct?

OpenStudy (anonymous):

because its already rooted?

OpenStudy (anonymous):

when you square it, the square root is gone

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you